# Magnetic Interaction in White Dwarf Binaries as Mechanism for Long-Period Radio Transients

### v1

@@ Abstract
A growing population of long-period radio transients has been discovered and their physical origin is still up to debate. Recently, a new such source named ILT J1101 + 5521 was discovered, which is in a white dwarf (WD) – M dwarf (MD) binary system, with the observed 125.5 min period being identified as the orbital period and the radio emission phase coinciding with the conjunction configuration when the MD is at the far end. We suggest that the radio emission properties of the system can be well explained within the framework of the unipolar inductor magnetic interaction model between the magnetized WD and the MD with low magnetization, with the electron cyclotron maser being the most likely radiation mechanism. This mechanism is similar to that of Jupiter decametric emission due to Jupiter-Io interaction. We suggest that this mechanism can interpret at least some long-period radio transients, especially the ultra-long period sub-population.
@@ 1 Introduction
Recently, a growing population of long-period radio transients (LPRTs) have been discovered (Hyman et al., 2005; Hurley-Walker et al., 2022, 2023, 2024; de Ruiter et al., 2024). These sources typically have a period from minutes to hours, usually with a duty cycle of the active window $\Delta t/P\ll 1$ (where $\Delta t$ is the active duration and $P$ is the period). The physical origin of these sources is unknown. The suggested engines include the spindown-powered magnetized white dwarf pulsars (Zhang & Gil, 2005; Katz, 2022), ultra-long-period magnetars (Wadiasingh et al., 2020), and regular magnetars with long-period precession (Zhu & Xu, 2006). A related white dwarf optical pulsar source with the similar period has an M dwarf companion Marsh et al. (2016) and may be powered by binary interactions (Geng et al., 2016). Multi-wavelength observations have given constraints on some of these models (Kaplan et al., 2008; Rea et al., 2022, 2024), but there has been no smoking gun observational clue that helps to nail down the physical origin of these systems.
@@
Recently, a newly discovered LPRT, named ILT J1101 + 5521, has been identified as a white dwarf (WD) – M dwarf (MD) system, and the measured $\sim$125.5-min period is identified as the orbital period (de Ruiter et al., 2024). Particularly, de Ruiter et al. (2024) found that the MD is behind the WD and in line with respect to Earth when the radio pulses are emitted. Incidentally, two other long period transients, i.e. the 3.56-hr-period optical white dwarf pulsar AR Scorpii (Marsh et al., 2016) and the 2.9-hr-period LPRT GLEAM-X J0704-37 (Hurley-Walker et al., 2024), are also associated with an MD companion. It seems that WD – MD binaries likely provide the preferred physical conditions to produce pulsed emission in such systems at least for these sources.
@@
In this paper, we propose that the coherent radio emission of ILT J1101 + 5521 is powered by a relativistic version of electron cyclotron maser mechanism in the magnetic field lines connecting the WD and the MD, with electromagnetic potential along the field lines excited through the unipolar inductor effect due to the orbital motion of the weakly magnetized MD moving in the magentosphere of the magneized WD. Such a mechanism has been successfully invoked to interpret the Jupiter decametric emission due to magnetic interactions between Jupiter and Io (Goldreich & Lynden-Bell, 1969), and it has been speculated to also operate in other systems as well, including ultra-compact double white dwarf binaries (Wu et al., 2002; Dall’Osso et al., 2006), short-orbital-period extrasolar super-earths (Laine & Lin, 2012), and binary neutron star or neutron star – black hole mergers (Hansen & Lyutikov, 2001; McWilliams & Levin, 2011; Lai, 2012; Piro, 2012; Wang et al., 2016). The paper is organized as follows. In Section 2, we discuss general constraints on physical models posted by the observations of ILT J1101 + 5521. In Section 3, we introduce the unipolar inductor magnetic internation model and apply it to ILT J1101 + 5521. In Section 4, we discuss various radiation mechanisms for ILT J1101 + 5521 and conclude that electron cylcotron maser emission is the most likely candidate. In Section 5 we discuss other LPRTs and suggest that the ultra-long-period sub-population of these sources may share the similar physical origin. The results are summarized in Section 6.
@@
Figure 1: The geometric sketch of the WD-MD unipolar induction magnetic interaction model. The left panel is the top view and the right panel is the side view. The blue and black companion denote the configurations with and without observed radio emission, respectively. The radio emission (purple wiggler) is produced by relativistic particles in the magnetic loop connecting the two stars. The line of sight (LOS) is denoted as red dashed line. The inclination angle $i$ is defined as the angle between the LOS and orbital plane normal vector. The separation distance between the WD and companion star is $a$.
2 Generic Observational Constraints
@@
(i) The observed period $\sim 125.5\ \rm min$ is consistent with the orbital period of the WD – MD binary system, with the radio pulses observed when the two stars are in conjunction and the MD is on the far side (de Ruiter et al., 2024). One can then estimate the semimajor axis of the binary as
@@
where $M=M_{\star}+M_{c}$ is the total mass of the binary system, and the characteristic values of $M_{\star}=0.6M_{\sun}$ and $M_{c}=0.2M_{\sun}$ are taken for the WD and the MD, respectively, following de Ruiter et al. (2024). The separation between the two stars ranges from $a(1-e)$ to $a(1+e)$, where $e$ is the eccentricity. The orbit is close to a circle for this source (de Ruiter et al., 2024), thus $e\simeq 0$.
@@
where $\Gamma$ and $\gamma$ are the bulk motion Lorentz factor and the typical electron Lorentz factor, respectively. For the binary system we are interested in, there is no bulk relativistic motion of the emitter, so we only normalize the electron Lorentz factor to 100. Since $T_{b}\gg T_{\rm b,incoh}$, the radio emission must be coherent.
@@
(iii) Since radio emission is only detected in certain orbital phases, the coherent radio emission must not be isotropically emitted. Geometric or relativistic beaming are needed (see Figure 1 for a possible geometry). The observed emission has high polarization, with the highest linear polarization degree $\sim 51\%$ for the brightest pulse and the circular polarization degree $<1.6\%$ and $<4.2\%$ for specific pulses (de Ruiter et al., 2024). These polarization properties are likely produced by relativistically moving electrons.
@@ 3.1 Physical picture
We consider a binary system consisting of a magnetized WD primary star with mass $M_{\star}$, radius $R_{\star}$, angular velocity $\Omega_{\star}$ and magnetic dipole moment $\mu=B_{\star}R_{\star}^{3}$. The companion MD star is non-magnetized with mass $M_{c}$ and radius $R_{c}$. The distance between the two stars is $a$. The angular velocity of the WD is $\Omega_{\star}=\xi\Omega$, where the orbital angular velocity is $\Omega=2\pi/P$ ($P$ is the orbital period).
@@
$\simeq(2.7\times 10^{-16}\ {\rm dynes})\ \left(\frac{M_{\star}}{0.6M_{\sun}}\right)\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{-2}.$
The electric force on one proton at the WD surface, on the other hand, may be calculated as
@@
$\simeq(9.3\times 10^{-8}\ {\rm dynes})\ \xi\left(\frac{P}{125.5\ \rm min}\right)^{-1}\left(\frac{R_{\star}}{0.01R_{\sun}}\right)\left(\frac{B_{\star}}{10^{7}\ \rm G}\right).$
@@
if $\xi\ll 10^{3}$. This suggests that the companion MD star is inside the magnetosphere of the WD.
We consider a toy model with $\bf\Omega_{\star}$ and $\bf\Omega$ aligned. The non-magnetized MD moving with respect to the WD magnetosphere would generate an electric potential drop of the magnitude of $\Phi\simeq 2R_{c}|\vec{E}|$ due to the unipolar induction effect. Here the factor of two is included since the circuit is repeated in both hemispheres. The induced electric field is
$\vec{E}=\frac{\vec{v}}{c}\times\vec{B}_{\rm bg},$ (7)
where $\vec{v}=|\vec{\Omega}_{\star}-\vec{\Omega}|\times\vec{r}=\Delta\Omega r\hat{\phi}$ is the companion star’s velocity relative to the WD magnetosphere. The background dipole magnetic field strength at the separation distance can be estimated as
$B_{\rm bg}=B_{\star}\left(\frac{R_{\star}}{a}\right)^{3}$ $\simeq(22\ {\rm G})\ \left(\frac{B_{\star}}{10^{7}\ \rm G}\right)\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{3}$ (8)
×$\times\left(\frac{M}{0.8M_{\sun}}\right)^{-1}\left(\frac{P}{125.5\ \rm min}\right)^{-2}.$
@@
$\Phi\simeq\frac{2\mu R_{c}}{ca^{2}}\Delta\Omega\simeq$ $(9.3\times 10^{8}\ {\rm statvolt})\ \zeta\left(\frac{B_{\star}}{10^{7}\ \rm G}\right)\left(\frac{R_{c}}{0.2R_{\sun}}\right)$ (9)
@@
where
Notice that the star can spin faster or slower than the orbit, i.e. $0<\xi\leq 1$ or $\xi>1$, but $\zeta>0$ is always satisfied because $\Delta\Omega$ is defined as the absolute value of $\Omega_{\star}$ and $\Omega$ difference. The current in the circuit may be calculated as $I={\Phi}/{2\mathcal{R}_{\rm mag}}$, where $\mathcal{R}_{\rm mag}={4\pi}/{c}$ is the resistance of the magnetosphere. One can then estimate the total electric power dissipation rate of the binary system as (Lai, 2012; Piro, 2012)
$\dot{E}$ =${}_{\rm diss}=\frac{2\Phi^{2}}{\mathcal{R}_{\rm mag}}\simeq\frac{4\mu^{2}R_{c}^{2}}{ca^{4}\pi}\Delta\Omega^{2}\simeq(2.1\times 10^{27}\ {\rm erg\ s^{-1}})\ \zeta^{2}$ (11)
×$\times\left(\frac{B_{\star}}{10^{7}\ {\rm G}}\right)^{2}\left(\frac{R_{c}}{0.2R_{\sun}}\right)^{2}\left(\frac{M}{0.8M_{\sun}}\right)^{-4/3}\left(\frac{P}{125.5\ \rm mins}\right)^{-14/3}.$
The highest observed specific radio luminosity is $\sim 7.8\times 10^{19}\ \rm erg\ s^{-1}\ Hz^{-1}$ (de Ruiter et al., 2024). One can see that the model can satisfy this constraint if $B_{\star}>2\times 10^{7}\ {\rm G}$ for $\zeta=1$. In reality, it is possible that $\Omega_{\star}>\Omega$ (or $\zeta>1$). This would increase the angular velocity difference which would further reduce the required $B_{\star}$.

### v2

# Magnetic Interactions in White Dwarf Binaries as Mechanism for Long-Period Radio Transients

@@ Abstract
A growing population of long-period radio transients has been discovered and their physical origin is still up to debate. Recently, a new such source named ILT J1101 + 5521 was discovered, which is in a white dwarf (WD) – red dwarf (RD) binary system, with the observed 125.5 min period in radio emission being identified as the orbital period and the radio emission occurs at the inferior conjunction of the WD. We suggest that the radio emission properties of the system can be well explained within the framework of the unipolar inductor magnetic interaction model between the magnetized WD and the RD with low magnetization, with a relativistic version of electron cyclotron maser emission being the most likely radiation mechanism. We suggest that this mechanism can interpret at least some long-period radio transients, especially the ultra-long period sub-population. Within this model, high energy emission in X-rays via relativistically boosted cyclotron radiation and $\gamma$-rays via inverse Compton scattering off stellar light are expected, but the predicted luminosities are relatively low. This model likely applies to the ultra-long period population of LPRTs. The short-period population of LPRTs is likely powered by other engines such as slow magnetars.
@@ 1 Introduction
Recently, a growing population of long-period radio transients (LPRTs) [^1] Also called “ultra-long-period objects (ULPOs) in the literature. have been discovered (Hyman et al., 2005; Hurley-Walker et al., 2022, 2023, 2024; de Ruiter et al., 2024). These sources typically have a period from minutes to hours, usually with a duty cycle of the active window $\Delta t/P\ll 1$ (where $\Delta t$ is the active duration and $P$ is the period). The physical origin of these sources is unknown. The suggested engines include spindown-powered magnetized white dwarf pulsars (Zhang & Gil, 2005; Katz, 2022), ultra-long-period magnetars (Wadiasingh et al., 2020; Cooper & Wadiasingh, 2024), and regular magnetars with long-period precession (Zhu & Xu, 2006). A related white dwarf optical pulsar source with a similar period has a red dwarf (M dwarf) companion (Marsh et al., 2016; Buckley et al., 2017) and may be powered by binary interactions (Geng et al., 2016). Multi-wavelength observations have given constraints on some of these models (Kaplan et al., 2008; Rea et al., 2022, 2024), but there has been no smoking gun observational clue that helps to nail down the physical origin of these systems.
@@
Recently, a newly discovered LPRT, named ILT J1101 + 5521, has been identified as a white dwarf (WD) – red dwarf (RD) system [^2] More specifically, the companion is identified as an M dwarf, but red dwarfs in general also include K dwarfs., and the measured $\sim$125.5-min period is identified as the orbital period (de Ruiter et al., 2024). Particularly, de Ruiter et al. (2024) found that the RD is behind the WD and in line with respect to Earth when the radio pulses are emitted. Incidentally, two other long period optical and radio transients, i.e. the 3.56-hr-period optical white dwarf pulsar AR Scorpii (Marsh et al., 2016; Buckley et al., 2017) and the 2.9-hr-period LPRT GLEAM-X J0704-37 (Hurley-Walker et al., 2024), are also found to be associated with an RD companion. It seems that WD – RD binaries likely provide the preferred physical conditions to produce pulsed emission in such systems at least for these sources.
@@
In this paper, we propose that the coherent radio emission of ILT J1101 + 5521 is powered by a relativistic version of electron cyclotron maser emission (ECME) mechanism in the magnetic field lines connecting the WD and the RD, with electromagnetic potential along the field lines excited through the unipolar inductor effect due to the orbital motion of the weakly magnetized RD moving in the magnetosphere of the magnetized WD. Such a mechanism has been successfully invoked to interpret the Jupiter decametric emission due to magnetic interactions between Jupiter and Io (Goldreich & Lynden-Bell, 1969), and it has been speculated to also operate in other systems as well, including ultra-compact double white dwarf binaries (Wu et al., 2002; Dall’Osso et al., 2006), short-orbital-period extrasolar super-earths (Laine & Lin, 2012), and binary neutron star or neutron star – black hole mergers (Hansen & Lyutikov, 2001; McWilliams & Levin, 2011; Lai, 2012; Piro, 2012; Wang et al., 2016). In particular, Chanmugam & Dulk (1982) and Dulk et al. (1983) have proposed that magnetic interactions in WD-RD systems can explain the radio emission of the magnetic cataclysmic variable, AM Her.
This paper is organized as follows. In Section 2, we discuss general constraints on physical models posted by the observations of ILT J1101 + 5521 and disfavor several radiation models. In Section 3, we introduce the unipolar inductor magnetic interaction model and apply it to ILT J1101 + 5521. In Section 4, we propose a relativistic version of ECME as the most likely radiation mechanism and discuss possible high-energy emission within this model. In Section 5 we discuss other LPRTs and suggest that the ultra-long-period sub-population of these sources may share similar physical origins. The results are summarized in Section 6.
@@
Figure 1: The geometric sketch of the WD-MD unipolar induction magnetic interaction model. The left panel is the top view and the right panel is the side view. The blue and black companion denote the configurations with and without observed radio emission, respectively. The radio emission (blue wiggler) is produced by relativistic particles in the magnetic loop connecting the two stars through relativistic ECME. High energy emissions are predicted, but are beamed in different directions, as marked as light blue wiggler (weak X-ray emission) via inverse Compton scattering, purple wiggler (strong X-ray emission) via relativistic boosting cyclotron radiation and red wiggler ($\gamma$-ray emission) via inverse Compton scattering off stellar light. The line of sight (LOS) is denoted as red dashed line. The inclination angle $i$ is defined as the angle between the LOS of radio emission and orbital plane normal vector. The separation distance between the WD and companion star is $a$.
2 Generic observational and theoretical constraints
@@
(i) The observed period $\sim 125.5\ \rm min$ is consistent with the orbital period of the WD – RD binary system, with the radio pulses observed when the binary is at inferior conjunction, i.e. the WD is between the Earth and the RD (de Ruiter et al., 2024). One can then estimate the semimajor axis of the binary as
@@
where $M=M_{\star}+M_{c}$ is the total mass of the binary system, and the characteristic values of $M_{\star}=0.6M_{\sun}$ and $M_{c}=0.2M_{\sun}$ are taken for the WD and the RD, respectively, following de Ruiter et al. (2024). The separation between the two stars ranges from $a(1-e)$ to $a(1+e)$, where $e$ is the eccentricity. The orbit is close to a circle for this source (de Ruiter et al., 2024), thus $e\simeq 0$ and the distance between the two stars is $a$.
@@
where $m_{e}$ is the electron mass, $\Gamma$ and $\gamma$ are the bulk motion Lorentz factor and the typical electron Lorentz factor, respectively. For the binary system we are interested in, there is no bulk relativistic motion of the emitter, so we only normalize the electron Lorentz factor to $100$. Since $T_{b}\gg T_{\rm b,incoh}$, the radio emission must be coherent.
@@
(iii) Since radio emission is only detected in certain orbital phases, the coherent radio emission must not be isotropically emitted. Geometric or relativistic beaming is needed (see Figure 1 for a possible geometry). The observed emission has high polarization, with the highest linear polarization degree $\sim 51\%$ for the brightest pulse and the circular polarization degree $<1.6\%$ and $<4.2\%$ for specific pulses (de Ruiter et al., 2024). These polarization properties are likely produced by relativistically moving electrons.
For completeness, we list other LPRTs detected so far in Table 1. One can see that besides ILT J1101 + 5521, a few other sources also have evidence of a possible RD companion, even though deep searches of optical counterparts of some others led to tight upper limits. X-ray emission was detected from AR Scorpii (Marsh et al., 2016; Buckley et al., 2017) and one recent source ASKAP J1832-0911 (Wang et al., 2024). This source was also detected by DART was named DART J1832-0911 (Li et al., 2024).
Table 1: Published LPRTs with distances, periods, flux density, luminosities, linear polarization degrees (in the radio band), possible optical counterparts and high energy (HE) counterparts. AR Scorpii is also included because of its possible similar origin. Corresponding references are listed below: [1]Hyman et al. (2005), [2]Kaplan et al. (2008), [3]Rea et al. (2022), [4]Hurley-Walker et al. (2022), [5]Hurley-Walker et al. (2023), [6]Men et al. (2025), [7]Caleb et al. (2024), [8]de Ruiter et al. (2024), [9]Hurley-Walker et al. (2024), [10]Li et al. (2024), [11]Wang et al. (2024), [12]Marsh et al. (2016), [13]Buckley et al. (2017), [14]Stanway et al. (2018), [15]Lee et al. (2025).
Source $D$ (kpc) $P\ ({\rm min})$ $S_{\nu}$ (Jy) $\Pi_{L}\ (\%)$ Optical counterparts HE counterparts Reference
GCRT J1745–3009 $<0.07$ 77 1 / No detection / [1],[2]
GLEAM-X J1627-52 $1.3\pm 0.5$ 18.18 5–40 $88\pm 1$ No detection / [3],[4]
GPM J1839-10 $5.7\pm 2.9$ 21 0.1–10 10–100 KD or MD? / [5],[6]
ASKAPJ1935+2148 4.85 53.8 0.119 $>90$ near IR source ? / [7]
ILT J1101 + 5521 0.504 125.5 0.041-0.256 $51\pm 6$ WD+RD / [8]
GLEAM-X J 0704-37 $1.5\pm 0.5$ 174 0.04 20–50 RD / [9]
DART/ASKAP J1832–0911 4.5 44 0.5-2 75-100 NS? X-ray [10],[11]
AR Scorpii $0.116\pm 0.016$ 213.6 0.003-0.012 $<1\%$ WD+RD X-ray [12],[13],[14]
ASKAP J1839-0756 $4.0\pm 1.2$ 387 0.022-1.8 90 No detection / [15]
In the following, we discuss some generic constraints on theoretical models based on observational data. The brightness temperatures of ILT J1101 + 5521 and other LPRTs are above the maximum allowed by incoherent emission, so the radio emission mechanism must be coherent. For ILT J1101 + 5521, we survey various emission mechanisms and can rule out the following mechanisms:
  • The required specific viewing geometry and the observed high linear polarization degrees suggest that the emitting electrons should be geometrically and relativistically beamed (otherwise, radio emission should be detected in all phases). This rules out the emission mechanisms invoking non-relativistic electrons, such as the non-relativistic version of ECME.
  • Type III solar flares emit coherent radio emission via collective Langmuir wave plasma emission (Melrose, 2017). For a static plasma, the required plasma number density is $\sim 10^{7}\ \rm cm^{-3}$ in order to have $\omega_{p}$ reach the observed frequency at $\sim 100$ MHz. This density is much greater than the Goldreich-Julian density of a WD magnetosphere (see Eq. (12). So at least the non-relativistic plasma emission is not suitable to explain the observations. More complicated relativistic plasma emission may be possible, but it is unclear how an instability can be driven from such a system.
  • Relativistic electrons can radiate in the background magnetic field via synchrotron radiation or curvature radiation. For typical parameters, the emission frequencies of synchrotron radiation and curvature radiation greatly exceed the typical observed radio frequency in LPRTs. For curvature radiation, one needs to lose the perpendicular momentum rapidly. In the characteristic emission region of interest in this paper, the background magnetic field $B_{\rm bg}$ field is not strong enough, and synchrotron cooling may not be sufficiently strong to make curvature radiation relevant. Also the traditional difficulties regarding formation and maintenance of the bunches also apply here. We therefore conclude that both synchrotron and curvature radiation are likely not the radiation mechanisms of these systems.
In the following, we will discuss a model to interpret the LPRT radio emission. The model includes two parts. The first part (Section 3 discusses a well-known unipolar induction model for magnetic interaction between the two stars. The second part (Section 4) proposes a relativistic version of the ECME mechanism to account for observed coherent radio emission from ILT J1101 + 5521 and other systems with similar properties of radio emission.
@@ 3.1 Physical picture
We consider a binary system consisting of a magnetized WD primary star with mass $M_{\star}$, radius $R_{\star}$, angular velocity $\Omega_{\star}$ and magnetic dipole moment $\mu=B_{\star}R_{\star}^{3}$, and a weakly magnetized RD with surface magnetic field strength $\leq{\rm kG}$, with mass $M_{c}$ and radius $R_{c}$. The distance between the two stars is $a$. The angular velocity of the WD is $\Omega_{\star}=\xi\Omega$, where the orbital angular velocity is $\Omega=2\pi/P$ ($P$ is the orbital period) and $\xi\equiv\Omega_{\star}/\Omega$ is a factor describing the ratio between the two angular velocities.
@@
$\simeq(2.7\times 10^{-16}\ {\rm dynes})\ \left(\frac{M_{\star}}{0.6M_{\sun}}\right)\left(\frac{R_{\star}}{0.01R_{\sun}}\right)^{-2},$
where $m_{p}$ is the proton mass. The electric force, on the other hand, may be calculated as
@@
$\simeq(9.3\times 10^{-9}\ {\rm dynes})\ \xi\left(\frac{P}{125.5\ \rm min}\right)^{-1}\left(\frac{R_{\star}}{0.01R_{\sun}}\right)\left(\frac{B_{\star}}{10^{6}\ \rm G}\right).$
@@
if $\xi\ll 10^{3}$. This suggests that the companion MD star is most likely inside the magnetosphere of the WD. The surface magnetic field of a typical RD is several kG from radio observations of magnetic cataclysmic variables (MCVs) (Barrett et al., 2020). For an RD with radius $R_{c}=0.2R_{\sun}$, the magnetic field strengths from the two stars are equal to $\sim 18\ {\rm G}$ at $r_{c}\simeq 2.67\times 10^{10}\ \rm cm$ away from the WD, which is close to the separation distance ($a\simeq 5.3\times 10^{10}\rm cm$). As long as the radio emission region is at $r<r_{c}$ from the WD (which is the case for the geometric configurations we consider), the WD magnetic field is the dominant component in the magnetosphere.
We consider a toy model with $\bf\Omega_{\star}$ and $\bf\Omega$ aligned. The MD moving with respect to the WD magnetosphere would generate an electric potential drop of the magnitude of $\Phi\simeq 2R_{c}|\vec{E}|$ due to the unipolar induction effect. Here the factor of two is included since the circuit is repeated in both hemispheres. The induced electric field is
$\vec{E}=\frac{\vec{v}}{c}\times\vec{B}_{\rm WD},$ (7)
where $\vec{v}=|\vec{\Omega}_{\star}-\vec{\Omega}|\times\vec{r}=\Delta\Omega r\hat{\phi}$ is the companion star’s velocity relative to the WD magnetosphere, $\hat{\phi}$ is the unit vector along azimuthal direction, $\vec{B}_{\rm WD}$ is the background magnetic field of the WD and
@@
$\Phi\simeq\frac{2\mu R_{c}}{ca^{2}}\Delta\Omega\simeq$ $(9.1\times 10^{7}\ {\rm statvolt})\ \zeta\left(\frac{B_{\star}}{10^{6}\ \rm G}\right)\left(\frac{R_{c}}{0.2R_{\sun}}\right)$ (9)
@@
The maximum Lorentz factor achievable by an electron accelerated in this potential can be estimated as
$\gamma_{\rm max}=\frac{q\Phi}{m_{e}c^{2}}$ $\simeq 5.4\times 10^{4}\ \zeta\left(\frac{B_{\star}}{10^{6}\ \rm G}\right)\left(\frac{R_{c}}{0.2R_{\sun}}\right)$ (10)
×$\times\left(\frac{M}{0.8M_{\sun}}\right)^{-2/3}\left(\frac{P}{125.5\ \rm min}\right)^{-7/3}.$
Notice that the star can spin faster than the orbit. Unless under the synchronous condition due to tidal lock (Wu & Wickramasinghe, 1993), $\zeta>0$ is usually satisfied because $\Delta\Omega$ is defined as the absolute value of $\Omega_{\star}$ and $\Omega$ difference. The current in the circuit may be calculated as $I={\Phi}/{2\mathcal{R}_{\rm mag}}$, where $\mathcal{R}_{\rm mag}={4\pi}/{c}$ is the resistance of the magnetosphere. One can then estimate the total electric power dissipation rate of the binary system as (Lai, 2012; Piro, 2012)
$\dot{E}$ =${}_{\rm diss}=\frac{2\Phi^{2}}{\mathcal{R}_{\rm mag}}\simeq(2.0\times 10^{25}\ {\rm erg\ s^{-1}})\ \zeta^{2}\left(\frac{B_{\star}}{10^{6}\ {\rm G}}\right)^{2}$ (11)
×$\times\left(\frac{R_{c}}{0.2R_{\sun}}\right)^{2}\left(\frac{M}{0.8M_{\sun}}\right)^{-4/3}\left(\frac{P}{125.5\ \rm mins}\right)^{-14/3}.$
The highest observed specific radio luminosity is $\sim 7.8\times 10^{19}\ \rm erg\ s^{-1}\ Hz^{-1}$ (de Ruiter et al., 2024). One can see that the model can satisfy this constraint if $\zeta>10$.
Chanmugam & Dulk (1982) and Dulk et al. (1983) assumed a near synchronous configuration ($\zeta\ll 1$) in their model to interpret radio emission of the magnetic cataclysmic variable, AM Her. Their typical Lorentz factor is $\sim 2$ because of a much smaller electric potential. In our model, we assume that the two stars are far from tidally locked and allow $\zeta$ as a free parameter. As long as $\zeta\gg 0$, electrons can be highly relativistic. It is worth mentioning that magnetized cataclysmic variables (mCVs), which are known sources of WD-RD systems with accretion onto the surface of WDs, typically have either $\zeta\gg 1$ (for the so-called intermediate-polars) when accretion is significant, or $\zeta\sim 0$ (for the so-called polars), when accretion does not exist and the system undergoes tidal locking (Warner, 1995). The coherent radio emission from LPRTs suggests that these sources should not have accretion (otherwise coherent emission would be quenched, as is the case of radio pulsars, Papitto et al. 2013). Indeed, these sources do not show observational signatures of known mCVs. They may be in the short transition phase from intermediate polars to synchronized systems, allowing the $\zeta$ value varying in a wide range between 0 and 10. In the rest of the discussion, we normalize $\zeta$ to unity but keep in mind that it could be an arbitrary value in a wide range. The unipolar mechanism we propose applies to a wide range of $\zeta$ values.